I have trouble simplifying radicals like this one : the square root of five over 12.

please help

Answer by stanbon(75887)

the square root of five over 12.
--------------------
sqrt(5/12)
---
= sqrt(5*12/12*12)
---
= sqrt(60/12^2)
---
= [sqrt(60)]/sqrt(12^2)
---
= [2sqrt(15)]/12
---
= (2/12)sqrt(15)
= (1/6)sqrt(15)
=======================
This is not the quickest way to do it but it is
the standard way.
=======================
Cheers,
stan H.